evaluate log 2(y -3) + log 4 base 2 =2
Answer by lwsshak3(11628)

evaluate log 2(y -3) + log 4 base 2 =2
log2(y-3) + log2(4) =2
log2[(y-3)*4]=log2(4)
4(y-3)=4
y-3=1
y=4